OpenStudy (anonymous):

Juanita and kauren are painting a circular mural on one wall of the community center . The area of the mural is 66 square meters . what is the radius (remember , pie=22/7 , a=pie r squared

OpenStudy (anonymous):

@Somy

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@Abhisar

OpenStudy (somy):

its circle right?

OpenStudy (anonymous):

yes I was thinking x=5

OpenStudy (somy):

formula for area of circle is Area= pi* r^2

OpenStudy (somy):

area is given and pi you know so find r

OpenStudy (anonymous):

oh kk

OpenStudy (somy):

66= (22/7) * r^2 so r^2= 66/ (22/7) so r= square root of the answer you get

OpenStudy (anonymous):

thnx

OpenStudy (somy):

u r welcome :))